A mixture Of milk ND Water 60lt which contains 10% water. How much water should be added so water May be 25%

Grade:12th pass

1 Answers

Arun
5 years ago

10% of 60 litres is 6 litres.
To make it 25% water, you'll have to add 'x' litres of water to the mixture.

(x+6)/(x+60) = 25/100
4x+24 = (x+60)
3x = 36
x = 12

So, if you add 12 litres of water to the mixture, you'll have 25% of water in the 72 litres of mixture, i.e., 18 litres.